OpenStudy (anonymous):

Radicle 250m^3 over 2n *m and n are different, it is not a mistake.

OpenStudy (anonymous):

are they both on sqrt like this: ?\[\sqrt{250m^{3}/2n}\]

OpenStudy (anonymous):

yeah, they're both radicles, except ones on top and others on bottom, not one big equation.

OpenStudy (anonymous):

And you're asked to?!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

asked to what

OpenStudy (anonymous):

\[\sqrt{250m^{3}} / \sqrt{2n}\]\[\sqrt{250m^{3}} / \sqrt{2n} * \sqrt{2n}/\sqrt{2n} = \sqrt{250m^{3}}\sqrt{2n} / 2n\]\[\sqrt{250m^{3}}\sqrt{2n} / 2n = \sqrt {500m^{3}n} \]

OpenStudy (anonymous):

thx